OpenStudy (hansc):

Given: ΔABC is a right triangle. Prove: a2 + b2 = c2 The following two-column proof with missing justifications proves the Pythagorean Theorem using similar triangles: Which is not a justification for the proof? Pieces of Right Triangles Similarity Theorem Side-Side-Side Similarity Theorem Substitution

OpenStudy (hansc):

Statement Justification Draw an altitude from point C to Line segment AB Let segment BC = a segment CA = b segment AB = c segment CD = h segment DB = y segment AD = x y + x = c c over a equals a over y and c over b equals b over x a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2

OpenStudy (greatlife44):

i think that formula is based off of the fact that you need to have a right triangle

OpenStudy (hansc):

could u help me?

OpenStudy (greatlife44):

@ganeshie8

OpenStudy (hansc):

could u help me? @ganeshie8